### v2

@@ 1.1 Overview of the paper
Our main findings provide several TTC characterizations through our fairness notions. Specifically, we show that: (1) along with strategy-proofness and pair efficiency, internal fairness and external fairness characterize TTC (Theorem 1); and (2) together with strategy-proofness, procedural fairness also characterize TTC (Theorem 2). Our results further highlight the exceptional performance of TTC from a fairness perspective.
Moreover, we consider an additional voluntary participation condition, the center lower bound, adapted from Chen and Feng (2025), which is weaker than Ekici (2024b)’s queuewise rationality. We also provide another TTC characterization by replacing external fairness with the center lower bound (Theorem 3).
Finally, we also explore “(strict) core” solution in our context. We modify the standard blocking notion to fit our model and correspondingly have a definition of our core solution, referred to as proper core, and show that it is single-valued and coincides with TTC (Theorem 4).
@@ 2 Related literature
More broadly, our paper contributes to the topic of allocation problems with incomplete priorities, e.g., Dur and Ünver (2019). In this type of problems, objects only have rank over partial agents and thus their priorities may be incomplete. Given this restriction, how to define fairness is crucial. In Dur and Ünver (2019), their fairness notion, respect for internal priorities is related to objects’ eligibility quotas as in their model, objects’ capacities may vary. However, in our setting, we cannot change the number of objects due to the constraints in reality, e.g., the supply of kidneys are scarce and may be fixed through a period. Instead, we focus on fairness within a framework where the size of the problem is fixed.
@@ 3.1 Multi-center allocation problems
The basic model setup is same to Ekici (2024b). There is a finite set of centers, $C$. Each center $c\in C$ consists of a finite set of agents (e.g., patients or workers) $N_{c}$ and a finite set of heterogeneous objects (e.g., kidneys or job positions) $O_{c}$, with $|N_{c}|=|O_{c}|$. [^2] In this paper, we do not consider the outside option because we want to keep balance in our model, we further discuss this point in Section 7. Each center $c$ has a strict priority $\succ_{c}$ over $N_{c}$. Each center’s priority of its members reflects some form of justice—whether based on need, queue time, or other criteria. However, the design of such priorities is beyond the scope of this paper, and we therefore assume that the priority structure is exogenously given.
@@ 3.2 Mechanisms and their properties
Originally, pair efficiency is viewed as an efficiency notion, which is weaker than Pareto efficiency. [^6] An allocation $x$ is Pareto efficient if there is no distinct allocation $y\neq x$ such that for each $i\in N$, $y_{i}\mathbin{R_{i}}x_{i}$. Here, one may reinterpret it as a fairness notion as it rules out the situation where two agents envy each other.
@@
By internal fairness, we rule out envy among agents within the same center based on that center’s priority. However, envy may still arise between agents from different centers, where no priority can be directly applied to determine whether such envy is “justified.” Nonetheless, as pointed out by Cheng et al. (2023), there may be an indirect relationship derived from priorities that can be applied to eliminate such envy. To address this, we introduce our second fairness property, external fairness, which is a modification of Morrill’s justness (Morrill, 2015). The intuitive interpretation of external fairness is as follows. If $i$ envies another agent $j$ from a different center and raises a complaint about the selected allocation, we must determine whether the complaint is justifiable. [^9] Note that here we use the term “complaint,” which differs from the term “violation of priority” we used earlier to define internal fairness. The fundamental distinction between the two arises from the difficulty we face: in the case of a priority violation, we can directly use a center’s priority to justify an agent’s envy. In contrast, for a complaint, no such priority can be directly applied. Thus, violations of priorities do not exist, making it harder to determine whether the envy is justified. Since $i$ and $j$ belong to different centers, no exogenously given priority can be directly applied to judge $i$’s complaint. However, as Morrill pointed out, we can consider a hypothetical scenario: If we accept this complaint and reassign $j$’s allotment, what would be the consequence? Specifically, would another agent $k$, who belongs to the same center as $i$ and has a higher priority, be harmed by this reassignment? If so, agent $k$, with a higher priority, may veto $i$’s complaint. Therefore, we should reject $i$’s complaint, as we prioritize $k$’s welfare over $i$’s, according to the priority of the center they both belong to.
@@
    $i$ and $j$ are from different centers, i.e., $c(i)\neq c(j)$; [^10] Similar to weak internal fairness (see Remark 1), here we can additionally require that $i$ and $f_{j}(R)$ belong to the same center to make sure that $i$ has a right to claim $f_{j}(R)$, this weaker version will not affect our results. and
@@
For each $c\in C$, let $i_{c}^{\tau}\equiv\max_{\succ_{c}}N_{c}^{\tau}$ be the agent with the highest priority among the remaining agents in $N_{c}^{\tau}$. For each $c\in C$, let $i_{c}^{\tau}$ point to his best preferred object among remaining objects according to $R_{i_{c}^{\tau}}$; and let each remaining object in $O_{c}^{\tau}$ point to $i_{c}^{\tau}$. There exists at least one cycle. For each cycle, define the trading opportunity set $T_{i}(R)$ for each involved agent $i$ as the set of objects pointing to $i$. Remove all agents and objects involved in cycles. Let $N^{\tau+1},O^{\tau+1}$ be the remaining agents and objects, respectively. If $N^{\tau+1}$ is not empty, then proceed to step $\tau+1$.
@@
Let $x\in X$ be the selected allocation. First, direct consumption. Suppose that an agent $i$ is unsatisfied at $x$ because he envies another agent $j$’s allotment $x_{j}$. If $x_{j}\in T_{i}(R)$, then it means that $i$ can directly consume $x_{j}$ from his trading opportunity set. Second, trading. Suppose that a set of agents $S$ who from different centers realize they can improve their welfare by trading objects from their respective trading opportunity sets. In this case, all agents in $S$ can claim $x$, the objects intended for trade, from their centers and proceed with the exchange: When cooperation across multiple centers is allowed, such trading represents the most easily achievable form of exchange.
Formally, an allocation $x$ is unfairly produced at $R$ if there exist a list of centers $c_{1},\ldots,c_{L}\in C$, a list of agents $(i_{1},\ldots,i_{L})\in\Pi_{\ell=1}^{L}N_{c_{\ell}}$, and a list of objects $(o_{1},\ldots,o_{L})\in\Pi_{\ell=1}^{L}T_{i_{\ell}}(R)$ such that for each $i\in\{i_{1},\ldots,i_{L}\}$, $o_{i+1}\mathbin{P_{i}}x_{i}$ (mod $L$). Otherwise $x$ is fairly produced at $R$. Note that if there is only one center (so $L=1$), then the condition represents the direct consumption, otherwise it represents the trading among agents $i_{1},\ldots,i_{L}$.
@@ 3.3 Top Trading Cycles
Each remaining agent points to his most-preferred remaining object given $R$. Each remaining object points to the agent with the highest priority among the remaining agents, according to $\succ_{O}$. There exists at least one cycle. Execute all cycles by assigning each agent involved in a cycle the object to which he points. Remove all agents and objects involved in a cycle. If some objects remain, then proceed to step $\tau+1$.
@@ 4.1 Internal fairness and external fairness
Note that from the proof of Theorem 1 below, one can see that our characterization in Theorem 1 also holds if we weaken internal fairness to weak internal fairness.
@@ 4.1.1 Proof of Theorem 1
Case One: $c(j)=c(i)$. In other words, $j\in N_{c(i)}$. By the selection of $R$, $f(R)$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. By the definition of TTC, each $k\in N_{c(i)}$ with higher priority than $i$ receives $x_{k}$ from some cycle in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Thus, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R)=x_{k}$. Therefore, $i\mathbin{\succ_{c(i)}}j$. So, weak internal fairness is violated.
Case Two: $c(j)\neq c(i)$, and hence $j\not\in N_{c(i)}$. By the selection of $R$, for each $R^{\prime}_{j}\in\mathcal{R}$, $\rho(R^{\prime}_{j},R_{-j})\geq t=\rho(R)$. Thus, $f(R^{\prime}_{j},R_{-j})$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})$. Recall that $j$ does not receive $x_{j}$ before step $t_{1}$ under $TTC(R)$. Thus, by the definition of TTC, $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})=\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Hence, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R^{\prime}_{j},R_{-j})=x_{k}$. Thus, $k$ does not depend on $j$. Also recall that $i\in N_{c(i)}$, $j\not\in N_{c(i)}$, and $f_{j}(R)=x_{i}\in O_{c(i)}$. Thus, $i$ externally justified envies $j$ at $f(R)$, contradicts external fairness of $f$. ∎
Thus, $|S_{\mathcal{C}}|=t_{2}=K\geq 2$. Without loss of generality, let $\mathcal{C}=(i_{1}\to x_{i_{1}}\to\ldots\to i_{K}\to x_{i_{K}}\to i_{1})$, and assume that $f_{i_{1}}(R)\neq x_{i_{1}}$. Let $c\equiv c(i_{1})$. That is, $i_{1}\in N_{c}$ and hence $x_{i_{K}}\in O_{c}$. Also, since $K\geq 2$, $x_{i_{1}}\not\in O_{c}$.
@@ 4.2.1 Proof of Theorem 2
Since $i\in I_{TTC}(R)$, $\mathcal{C}$ is not executed at $f(R)$. Hence, Lemma 2 and procedural fairness of $TTC$ together imply that $f_{i}(R)=e_{i}(R)$. This implies that the agent who points to $e_{i}(R)$ in $\mathcal{C}$, say agent $j$, does not receive $e_{i}(R)$ at $f(R)$. By applying similar arguments to agent $j$, we see that $f_{j}(R)=e_{j}(R)$. By repeating this argument for all agents in $S_{\mathcal{C}}$, we conclude that for each agent $k\in S_{\mathcal{C}}$, $f_{k}(R)=e_{k}(R)$. Consequently, $S_{\mathcal{C}}\subseteq I_{TTC}(R)$.
@@ 5 Relation to queuewise rationality
: Let $\delta=(1,2,3)$ be an order of $N$. Let $R\in\mathcal{R}^{N}$. Let $Y(0)\subseteq X$ be the set of allocations that satisfy queuewise rationality.
@@
Example 3 (TTC with artificial endowments).
Let $e(R)\in X$ be such that for each $i\in N$, $e_{i}(R)=o_{i}(R)$. Let $e(R)$ be an “artificial” endowment allocation. We run the TTC algorithm with respect to $e(R)$, i.e., let each $e_{i}(R)$ points to $i$. Define the corresponding mechanism as $TTC^{e}$. This mechanism satisfies queuewise rationality and external fairness, but violates internal fairness. $\diamond$
@@ 6.1 Core-stability
Still, the first condition can be directly applied to our context. However, the second condition needs to be modified because in our model, an agent may not own a single object, but rather a set of objects associated with the center to which he belongs. Thus, it is natural to require that what agents can bring to form a blocking based on the underlying priority structure. [^17] Aligning with this, Ekici (2024b) modifies the self-autarky condition in a different way: he requires that for each $i\in S$, all agents who are from the same center and have higher priorities than $i$ must also be in $S$. Aligning with our idea of procedural fairness, here we adjust the self-autarky condition by requiring that each member brings one object from his trading opportunity set.

### v3

@@ 1.1 Overview of the paper
Our main findings provide several TTC characterizations through our fairness notions. Specifically, we show that: (1) along with strategy-proofness and pair efficiency, internal fairness and external fairness characterize TTC (Theorem 1); and (2) together with strategy-proofness, procedural fairness also characterize TTC (Theorem 2). Our results further highlight the exceptional performance of TTC from a fairness perspective. Moreover, we consider an additional voluntary participation condition, the center lower bound, adapted from Chen and Feng (2025), which is weaker than Ekici (2024b)’s queuewise rationality. We also provide another TTC characterization by replacing external fairness with the center lower bound (Theorem 3). Finally, we also explore “(strict) core” solution in our context. We modify the standard blocking notion to fit our model and correspondingly have a definition of our core solution, referred to as proper core, and show that it is single-valued and coincides with TTC (Theorem 4).
@@ 2 Related literature
More broadly, our paper contributes to the topic of allocation problems with incomplete priorities, e.g., Dur and Ünver (2019). In this type of problems, objects only have rank over partial agents and thus their priorities may be incomplete. Given this restriction, how to define fairness is crucial. In Dur and Ünver (2019), their fairness notion, respect for internal priorities is related to objects’ eligibility quotas as in their model, objects’ capacities may vary. However, in our setting, we cannot change the number of objects due to the constraints in reality, e.g., the supply of organs are scarce and may be fixed through a period. Instead, we focus on fairness within a framework where the size of the problem is fixed.
@@ 3.1 Multi-center allocation problems
The basic model setup is same to Ekici (2024b). There is a finite set of centers, $C$. Each center $c\in C$ consists of a finite set of agents (e.g., patients or workers) $N_{c}$ and a finite set of heterogeneous objects (e.g., organs or job positions) $O_{c}$, with $|N_{c}|=|O_{c}|$. [^2] In this paper, we do not consider the outside option because we want to keep balance in our model, we further discuss this point in Section 7. Each center $c$ has a strict priority $\succ_{c}$ over $N_{c}$. Each center’s priority of its members reflects some form of justice—whether based on need, queue time, or other criteria. However, the design of such priorities is beyond the scope of this paper, and we therefore assume that the priority structure is exogenously given.
@@ 3.2 Mechanisms and their properties
Originally, pair efficiency is viewed as an efficiency notion, which is weaker than Pareto efficiency. [^6] An allocation $x$ is Pareto efficient if there is no distinct allocation $y\neq x$ such that for each $i\in N$, $y_{i}\mathbin{R_{i}}x_{i}$. Here, one may reinterpret it as a fairness notion as it rules out the situation where two agents envy each other, aligning with the idea that swapping two objects eliminates envy (Echenique et al., 2025).
@@
By internal fairness, we rule out envy among agents within the same center based on that center’s priority. However, envy may still arise between agents from different centers, where no priority can be directly applied to determine whether such envy is “justified.” Nonetheless, as pointed out by Cheng et al. (2023), there may be an indirect relationship derived from priorities that can be applied to eliminate such envy. To address this, we introduce our second fairness property, external fairness, which is a modification of Morrill’s justness (Morrill, 2015). The intuitive interpretation of external fairness is as follows. If $i$ envies another agent $j$ from a different center and raises a complaint about the selected allocation, we must determine whether the complaint is justifiable. [^9] Note that here we use the term “complaint,” which differs from the term “violation of priority” we used earlier to define internal fairness. The fundamental distinction between the two arises from the difficulty we face: in the case of a priority violation, we can directly use a center’s priority to justify an agent’s envy. In contrast, for a complaint, no such priority can be directly applied. Thus, violations of priorities do not exist, making it harder to determine whether the envy is justified. Since $i$ and $j$ belong to different centers, no exogenously given priority can be directly applied to judge $i$’s complaint. However, as Morrill pointed out, we can consider a hypothetical scenario: If we accept this complaint and reassign $j$’s allotment, what would be the consequence? Specifically, would another agent $k$, who belongs to the same center as $i$ and has a higher priority, be directly harmed by this reassignment? [^10] Here, by “directly harmed” we mean that $k$’s allotment is directly reassigned to $j$, which aligns with the idea of Morrill (2015). In other words, we formalize harm by considering only immediate effects. More complex chains, in which $j$’s deviation indirectly worsens $k$’s assignment without $j$ ever receiving $k$’s allotment, are not considered, as the current notion is sufficient to formalize our fairness concept and guarantee positive results. If so, agent $k$, with a higher priority, may veto $i$’s complaint. Therefore, we should reject $i$’s complaint, as we prioritize $k$’s welfare over $i$’s, according to the priority of the center they both belong to.
@@
    $i$ and $f_{j}(R)$ belong to the same center, while $j$ does not, i.e., $c(i)\neq c(j)$ and $f_{j}(R)\in O_{c(i)}$; and
@@
Remark 2.
Strong external fairness
Unlike internal fairness and weak internal fairness (see Remark 1), here we cannot drop the requirement that $i$ and $f_{j}(R)$ belong to the same center. Otherwise, any cross-center envy that a highest-priority agent $i$ has toward any $j(\not\in N_{c(i)})$ would be classified as externally justified (see Footnote 11), leading to an impossibility result. In other words, no mechanism can satisfy this stronger version of external fairness. $\diamond$
@@
For each $c\in C$ with $N_{c}^{\tau}\neq\emptyset$, let $i_{c}^{\tau}\equiv\max_{\succ_{c}}N_{c}^{\tau}$ be the agent with the highest priority among the remaining agents in $N_{c}^{\tau}$. For each $c\in C$, let $i_{c}^{\tau}$ point to his best preferred object among remaining objects according to $R_{i_{c}^{\tau}}$; and let each remaining object in $O_{c}^{\tau}$ point to $i_{c}^{\tau}$. Define the trading opportunity set $T_{i_{c}^{\tau}}(R)$ as the set of objects pointing to $i_{c}^{\tau}$. Also, given agents and objects’ pointing graph, there exists at least one cycle. Remove all agents and objects involved in cycles. Let $N^{\tau+1},O^{\tau+1}$ be the remaining agents and objects, respectively. If $N^{\tau+1}$ is not empty, then proceed to step $\tau+1$.
@@
Let $x\in X$ be the selected allocation. First, direct consumption. Suppose that an agent $i$ is unsatisfied at $x$ because he envies another agent $j$’s allotment $x_{j}$. If $x_{j}\in T_{i}(R)$, then it means that $i$ can directly consume $x_{j}$ from his trading opportunity set. Second, trading. Suppose that a set of agents $S$ who from different centers realize they can improve their welfare by trading objects from their respective trading opportunity sets. In this case, all agents in $S$ can deviate from $x$, by using the objects from their trading opportunity sets and proceeding with the exchange. When cooperation across multiple centers is allowed, such trading represents the most easily achievable form of exchange.
Formally, an allocation $x$ is unfairly produced at $R$ if there exist a list of centers $c_{1},\ldots,c_{L}\in C$, a list of agents $(i_{1},\ldots,i_{L})\in\Pi_{\ell=1}^{L}N_{c_{\ell}}$, and a list of objects $(o_{1},\ldots,o_{L})\in\Pi_{\ell=1}^{L}T_{i_{\ell}}(R)$ such that for each $i_{\ell}\in\{i_{1},\ldots,i_{L}\}$, $o_{i_{\ell}+1}\mathbin{P_{i_{\ell}}}x_{i_{\ell}}$ (mod $L$). Otherwise $x$ is fairly produced at $R$. Note that if there is only one center (so $L=1$), then the condition represents the direct consumption, otherwise it represents the trading among agents $i_{1},\ldots,i_{L}$.
@@ 3.3 Top Trading Cycles
Each remaining agent points to his most-preferred remaining object given $R$. Each remaining object points to the agent with the highest priority among the remaining agents, according to $\succ_{O}$. [^16] Note that, since for each center $c$ the balance constraint $|N_{c}|=|O_{c}|$ holds, if an object in $O_{c}$ remains at step $t$, then some agent in $N_{C}$ must also remain at step $t$. As a result, each object in $O_{c}$ points only to agents in $N_{c}$. There exists at least one cycle. Execute all cycles by assigning each agent involved in a cycle the object to which he points. Remove all agents and objects involved in a cycle. If some objects remain, then proceed to step $\tau+1$.
@@ 4.1 Internal fairness and external fairness
Note that from the proof of Theorem 1 below, one can see that our characterization in Theorem 1 also holds if internal fairness is weakened to weak internal fairness (see Remark 1).
@@ 4.1.1 Proof of Theorem 1
Case One: $c(j)=c(i)$. In other words, $j\in N_{c(i)}$. By the selection of $R$, $f(R)$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. By the definition of TTC, each $k\in N_{c(i)}$ with higher priority than $i$ receives $x_{k}$ from some cycle in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Thus, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R)=x_{k}$. Therefore, $i\mathbin{\succ_{c(i)}}j$. So, (weak) internal fairness is violated.
Case Two: $c(j)\neq c(i)$, and hence $j\not\in N_{c(i)}$. By the selection of $R$, for each $R^{\prime}_{j}\in\mathcal{R}$, $\rho(R^{\prime}_{j},R_{-j})\geq t=\rho(R)$. Thus, $f(R^{\prime}_{j},R_{-j})$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})$. Recall that $j$ does not receive $x_{j}$ before step $t_{1}$ under $TTC(R)$. Thus, by the definition of TTC, $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})=\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Hence, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R^{\prime}_{j},R_{-j})=x_{k}$. Thus, $k$ does not depend on $j$. Also recall that $i\in N_{c(i)}$, $f_{j}(R)=x_{i}\in O_{c(i)}$, and $j\not\in N_{c(i)}$. Thus, $i$ externally justified envies $j$ at $f(R)$, contradicts external fairness of $f$. ∎
Note that our two fairness notions are only used to prove Claim 1.
By Claim 1, $|S_{\mathcal{C}}|=t_{2}=K\geq 2$. Without loss of generality, let $\mathcal{C}=(i_{1}\to x_{i_{1}}\to\ldots\to i_{K}\to x_{i_{K}}\to i_{1})$, and assume that $f_{i_{1}}(R)\neq x_{i_{1}}$. Let $c\equiv c(i_{1})$. That is, $i_{1}\in N_{c}$ and hence $x_{i_{K}}\in O_{c}$. Also, since $K\geq 2$, $x_{i_{1}}\not\in O_{c}$.
@@ 4.2.1 Proof of Theorem 2
Since $i\in I_{TTC}(R)$, $\mathcal{C}$ is not executed at $f(R)$. Thus, $f_{i}(R)\neq TTC_{i}(R)$. By (1), $f_{i}(R)\in\{e_{i}(R),TTC_{i}(R)\}$. Therefore, $f_{i}(R)=e_{i}(R)$. This implies that the agent who points to $e_{i}(R)$ in $\mathcal{C}$, say agent $j$, does not receive $e_{i}(R)$ at $f(R)$. By applying similar arguments to agent $j$, we see that $f_{j}(R)=e_{j}(R)$. By repeating this argument for all agents in $S_{\mathcal{C}}$, we conclude that for each agent $k\in S_{\mathcal{C}}$, $f_{k}(R)=e_{k}(R)$. Consequently, $S_{\mathcal{C}}\subseteq I_{TTC}(R)$.
@@ 5 Relation to queuewise rationality
: Let $\delta=(1,2,3)$ be an order of $N$. Let $R\in\mathcal{R}^{N}$. Let $Y(0)\subseteq X$ be the set of allocations that satisfy queuewise rationality at $R$.
@@
Example 3 (Two steps: SD then TTC).
Given $R$, for each center $c$, first run serial dictatorship over $O_{c}$ with respect to $\succ_{c}$: the highest-priority agent in $N_{c}$ picks up his most preferred object in $O_{c}$, the second-highest priority agent picks his most preferred object among the remaining objects in $O_{c}$, and so on. Denote the resulting allocation by $e(R)$. Then, let $e(R)$ be an “artificial” endowment allocation and run the TTC algorithm with respect to $e(R)$, i.e., let each $e_{i}(R)$ points to $i$. This two-step mechanism satisfies queuewise rationality and external fairness, but violates internal fairness. $\diamond$
@@ 6.1 Core-stability
Still, the first condition can be directly applied to our context. However, the second condition needs to be modified because in our model, an agent may not own a single object, but rather a set of objects associated with the center to which he belongs. Thus, it is natural to require that what agents can bring to form a blocking based on the underlying priority structure. Aligning with our idea of procedural fairness, here we adjust the self-autarky condition by requiring that each member brings one object from his trading opportunity set.
@@
Two points are worth noting about our self-autarky condition. First, in the special case where each center contains only one agent, it reduces to the classical self-autarky condition used in the definition of the core. Second, recall that an agent’s trading opportunity set depends on the preferences of agents with higher priority. This restriction implicitly implies that whether an agent can (or cannot) bring an object to form a blocking depends on the views of higher-priority agents. [^18] Aligning with this, Ekici (2024b) modifies the self-autarky condition in a different way: he requires that for each $i\in S$, all agents who are from the same center and have higher priorities than $i$ must also be in $S$.